Check, whether the collection of G, given by

G'= { ] 1/n+1,1/n [ : n∈N } is a open curve of ]0,1[


1
Expert's answer
2021-05-07T10:09:53-0400

"G={(5-n,5+\dfrac{1}{n})|n\in N}"


Let "B_n=(5-n,5+\dfrac{1}{n})"


Then,"B_1=(4,6) ,B_2=(3,5+\dfrac{1}{2})"

   "B_3=(2,5+\dfrac{1}{3}), B_4=(1,5+\dfrac{1}{4})"

   "B_5=(0,5+\dfrac{1}{5}),....."


Here "B_2 B_1,B_3 B_2,...., B_{n+1} B_n"

Hence the collections of sets {"B_n" } is notnested.


Also, we see that, {"B_n" } are not mutually disjoint a"B_1 \cup B_2\neq \phi, B_2\cap B_3\neq \phi,..., B_n\cap B_{n+1}\neq \phi"


Hence The given set G is not an open curve on [0,1]